OpenStudy (anonymous):

y=2* {(-1,__),(0,__),(1,__),(2,__),(3,__),(4,__)}

OpenStudy (anonymous):

\[y=2^x \] right? \[2^{-1}=\frac{1}{2}\] \[2^0=1\] \[2^1=2\] etc

OpenStudy (anonymous):

thank you :)